OpenStudy (anonymous):

Find the exact value sec^2pi/6 - 4

OpenStudy (anonymous):

sec(2pi/6)-4?

OpenStudy (anonymous):

sec^(2pi/6)-4 has no angle, it just tells you the power to raise sec to

OpenStudy (anonymous):

\[\sec ^2\pi/6-4\]

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

How'd you get that? @larryboxaplenty

OpenStudy (anonymous):

Is it sec^2(2pi/6) -4, or sec^2( (2pi/6) -4)?

OpenStudy (anonymous):

If the -4 is not in the sec^2, then:

OpenStudy (anonymous):

sec^2(x) = 1/cos(x) * 1/cos(x). 1/cos(pi/6) = 2/sqrt(3) 1/cos(pi/6) * (1/cos(pi/6) = (2/sqrt(3))^2 = 4/3 4/3 - 4 = 8/3